OpenStudy (anonymous):

Determine the nature of the solutions of the equation. x^2-14x+49=0 What does the equation have? Is my answer correct "one real solution"?

OpenStudy (anonymous):

tey only one real solution...

OpenStudy (anonymous):

x^2-14x+49=0 (x-7)^2=0 x=7

OpenStudy (anonymous):

just calculate b^2 - 4ac i.e 14^2 - 4(1)(49) if it is positive i.e greater than zero, you will have two distinct real roots if it is zero, you will have two real equal roots if it is negative i.e. less than zero, you will have imaginary roots here b^2 - 4ac = 196 - 196 = 0 , so it has tow real equal roots

OpenStudy (anonymous):

Hope this clears it for you..... If yes, pls click GOOD ANSWER button to show your satisfaction